# Spin pumping into quantum spin nematic states

### v1

@@ IV.1 Zero magnetic field
Figure 2: The increase of the Gilbert damping for the dirty interface as a function of the FMR frequency $\omega_{\bm{0}}$. The parameters are set as $J_{1}=0$, $J_{2}=-{\cal J}$, and $h=0$. The solid and dashed lines represent $\theta=0$ and $\theta=\pi/2$, respectively.
@@
The increase of the Gilbert damping significantly depends on whether the interface is clean or dirty. In the case of a clean interface, the Gilbert damping is not modulated, that is, $\delta\alpha_{\rm G}=0$ because the coefficient in Eq. (46) becomes zero. In contrast, for a dirty interface, $\delta\alpha_{\rm G}$ has a finite value. In Fig. 2, we plot $\delta\alpha_{\rm G}/\delta\alpha_{{\rm G},0}$ as a function of the FMR frequency $\omega_{\bm{0}}$, where $\delta\alpha_{{\rm G},0}=2\pi S_{0}|\bar{T}|^{2}{\cal A}/{\cal J}^{2}$ is a normalization factor. As shown in the figure, the enhancement of Gilbert damping for $\theta=\pi/2$ is twice that of $\theta=0$. In general, $\delta\alpha_{\rm G}$ is proportional to $1+\sin^{2}\theta$. This dependence arises from the fact that $\delta\alpha_{\rm G}$ is a sum of the two contributions from two types of Schwinger bosons; the contribution of the $z^{\prime}$-component Schwinger bosons is independent of $\theta$, while that of the $x^{\prime}$-component Schwinger bosons is proportional to $\sin^{2}\theta$ (see also Eqs. (43)-(45)).
IV.2 Dirty interface under a finite magnetic field
@@
Figure 3: The increase of the Gilbert damping for the dirty interface is shown as a function of the FMR frequency $\omega_{\bm{0}}$ for $J_{1}=0$ and $J_{2}=-{\cal J}$. (a) The magnetic-field dependence at $\theta=0$. The three curves correspond to $h={\cal J}$, $3{\cal J}$, and $5{\cal J}$, respectively. (b) The angle dependence at $h=3{\cal J}$. The red, blue, and green solid lines correspond to the contribution of $C_{z}$ for $\theta=0$, $\theta=\pi/2$, and $\theta=\pi$, respectively. The dashed line represents the contribution of $C_{\perp}$ at $\theta=\pi/2$. We note that $C_{\perp}(\theta=0)=C_{\perp}(\theta=\pi)=0$.
@@
Fig. 3(a) shows $\delta\alpha_{\rm G}/\delta\alpha_{{\rm G},0}$ as a function of the FMR frequency $\omega_{\bm{0}}$ in the presence of an external magnetic field. As the magnetic field increases, $\delta\alpha_{{\rm G}}(\omega_{\bm{0}})$ grows because the net magnetization in the spin-nematic insulator increases and the peak becomes sharper reflecting changes in the dispersion relation of the Schwinger bosons (see also Fig. 4(d)-(f)). Fig. 3(b) shows the increase of the Gilbert damping at $\theta=0$, $\theta=\pi/2$, $\theta=\pi$ with the magnetic field fixed at $h=3{\cal J}$, where the contributions of the two kinds of Schwinger bosons are plotted separately. We note that the contribution from the perpendicular Schwinger bosons vanishes for $\theta=0,\pi$. The peak at $\omega_{\bm{0}}=2.9{\cal J}$ is the contribution of the $z^{\prime}$-component Schwinger bosons, and the peak at $\omega_{\bm{0}}=3.2{\cal J}$ is that of the perpendicular component of the Schwinger bosons. As seen in the figure, the two components of the bosons have different dependencies on $\theta$.
Figure 4: The top row shows the increase of the Gilbert damping for the dirty interface as a function of the FMR frequency $\omega_{\bm{0}}$ and $\theta$. The bottom row shows the dispersion of the Schwinger bosons of $\varepsilon_{{\bm{k}}z}$ and $\varepsilon_{{\bm{k}}{\perp}}$. The parameters are set as $J_{1}=0$ and $J_{2}=-{\cal J}$. In each row, figures are arranged from left to right, $h={\cal J}$, $h=3{\cal J}$, and $h=5{\cal J}$.
@@
To clarify the dependence on the magnetic-field orientation, we show the increase of the Gilbert damping as a function of $\omega_{\bm{0}}$ and $\theta$ for $h={\cal J},3{\cal J}$, and $5{\cal J}$ in Fig. 4 (a), (b), and (c), respectively. The peak height at a finite frequency varies as the angle $\theta$ changes. The shape and position of the peak of $\delta\alpha_{\rm G}$ are determined by the corresponding dispersion of the two Schwinger bosons, $\varepsilon_{{\bm{k}}\nu}$, as shown in Fig. 4(d)-(f). As expected from the form of Eq. (43), the peak frequency of $\delta\alpha_{\rm G}$ corresponds to the energy of the flat part in the dispersion of the Schwinger bosons. The low-frequency shoulder structure of $\delta\alpha_{\rm G}$ reflects the low-energy states of the Schwinger bosons, though its weight is much weaker than that of the main peak.
@@
Let us examine the detailed features for each strength of the magnetic field. For $h={\cal J}$, the dispersions of the two Schwinger bosons are almost identical except for the $\Gamma$ point, as shown in Fig. 4(d). As a result, $\delta\alpha_{\rm G}$ has a single peak and becomes significant near $\theta=\pi/2,3\pi/2$ as shown in Fig. 4(a). For a higher magnetic field of $h=3{\cal J}$, a new high-energy peak appears as shown in Fig. 4(b) The low-energy peak around $\omega_{\bm{0}}=2.9{\cal J}$ becomes significant near $\theta=0$, while the high-energy peak around $\omega_{\bm{0}}=3.2{\cal J}$ appears around $\theta=\pi/2,3\pi/2$. This behavior reflects the dispersion relations of the two Schwinger bosons as shown in Fig. 4(e). The positions of the double peaks correspond to the flat bands, whose energies are different for the two types of bosons, while their weights have a different dependencies on $\theta$, as seen also in Eqs. (43)-(45). Fig. 4(c) shows the increase of the Gilbert damping for a higher magnetic field, $h=5{\cal J}$. Under a strong magnetic field, the contribution of the perpendicular component of the Schwinger bosons almost vanishes, while that of the $z^{\prime}$-component bosons becomes dominant. As a result, $\delta\alpha_{\rm G}$ has only a single low-energy peak due to the flat part in the dispersion of the $z^{\prime}$ Schwinger bosons.
IV.3 Clean interface under a finite magnetic field
Figure 5: The increase of the Gilbert damping for the clean interface as a function of the FMR frequency $\omega_{\bm{0}}$ for $J_{1}=0$, $J_{2}=-{\cal J}$. (a) The magnetic-field dependence for $\theta=0$. The three curves correspond to $h={\cal J}$, $3{\cal J}$, and $5{\cal J}$, respectively. (b) The angle dependence for $h=3{\cal J}$.
Next, we show the result for the clean interface. In Fig. 5(a), we plot $\delta\alpha_{\rm G}/\delta\alpha_{{\rm G},0}^{\prime}$ as a function of the FMR frequency $\omega_{\bm{0}}$, where $\delta\alpha_{{\rm G},0}^{\prime}=2\pi S_{0}|\bar{T}|^{2}/{\cal J}^{2}$. In contrast to the case of the dirty interface, the peak position depends on the magnetic field, reflecting the energy gap of the massive Schwinger boson, $\varepsilon_{{\bm{k}}z}$, at ${\bm{k}}={\bm{0}}$. In Fig. 5(b), we present the increase of the Gilbert damping for $h=3{\cal J}$ as a function of $\omega_{\bm{0}}$ and $\theta$. In the case of $\theta=\pi$, the Gilbert damping is not modulated, meaning $\delta\alpha_{\rm G}=0$. This is consistent with the fact that the coefficient becomes zero when $\theta=\pi$ and ${\bm{k}}=0$ in Eq. (46). On the other hand, the peak becomes significant near $\theta=0,2\pi$.
@@ V Summary
In this study, we examined spin pumping into a spin-nematic (SN) state in a junction system composed of a ferromagnetic insulator and a spin-nematic insulator. To clarify its qualitative features, we employed a spin-1 bilinear-biquadratic model to realize the spin-nematic phase and calculated the increase of the Gilbert damping using the mean-field theory based on the Schwinger boson formalism. We considered two types of interfaces—clean and dirty—using simple models with interfacial exchange coupling, both with and without momentum conservation. In the absence of a magnetic field, spin pumping is completely suppressed for a clean interface. However, for a dirty interface, the increase of the Gilbert damping exhibits a peak structure with a low-energy shoulder as a function of the FMR frequency, reflecting the energy dispersion of the two Schwinger bosons. The orientation angle of boson condensation also influences the increase of the Gilbert damping. In the presence of the magnetic field, the different angular dependence of the two kinds of bosons becomes more pronounced for a dirty interface, and the peak structure of the increase of Gilbert damping exhibits a complex dependence on the magnetic field orientation. For a clean interface, on the other hand, the increase of the Gilbert damping is governed by the energy gap of one of the two Schwinger boson modes at zero wavenumber and exhibits a single peak with strong dependence on the magnetic field orientation. Our results will be helpful in identifying the spin-nematic phase through spin pumping, a widely used technique in spintronics research.
@@ Acknowledgements
The authors thank M. Matsuo for his helpful discussions. T. I. is supported by the International Graduate Program of Innovation for Intelligent World (IIW) of the University of Tokyo and the International Program of Institute for Solid State Physics, the University of Tokyo. T. K. acknowledges the support of the Japan Society for the Promotion of Science (JSPS KAKENHI Grant No. JP20K03831).
@@ Appendix A Detailed calculation of the spin-spin correlation function
$S^{+^{\prime}}_{\bm{k}}$ $=i\sqrt{M}\,\bigl{[}(C^{-}_{\mu}c^{z}_{{\bm{k}}}-C^{+}_{\mu}s^{z}_{{\bm{k}}})\alpha_{{\bm{k}}z}^{\dagger}+(C^{-}_{\mu}s^{z}_{{\bm{k}}}-C^{+}_{\mu}c^{z}_{{\bm{k}}})\alpha_{-{\bm{k}}z})\bigr{]},$ (50)
@@
$\chi^{z^{\prime}z^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)$ $=M\cos^{2}\mu(c_{\bm{k}}^{\perp}-s_{\bm{k}}^{\perp})^{2}\bigl{[}G_{{\bm{k}}z}(\tau)+G_{{\bm{k}}z}(-\tau)\bigr{]},$ (55)
@@
Using the diagonalized Hamiltonian, Eq. (22), this propagator is calculated as

### v2

@@ IV.1 Zero magnetic field
Figure 3: The increase of the Gilbert damping as a function of the FMR frequency $\omega_{\bm{0}}$ for (a) the SN insulator and (b) the CAF insulator. The solid and dashed lines represent $\theta=0$ and $\theta=\pi/2$, respectively. The parameters are set as $J_{1}=0$, $J_{2}=-{\cal J}$, $J_{3}={\cal J}$ and $h=0$. We note that the Gilbert damping is independent of $\theta$ for the CAF insulator.
@@
In Fig. 3 (a) and (b), we plot $\delta\alpha_{\rm G}/\delta\alpha_{{\rm G},0}$ as a function of the FMR frequency $\omega_{\bm{0}}$ for the SN insulator and the CAF insulator, respectively. Here, $\delta\alpha_{{\rm G},0}=2\pi S_{0}|\bar{T}|^{2}/{\cal J}^{2}$ is a normalization factor. We note that in the absence of the magnetic field ($h=0$), the model for the CAF insulator leads to collinear AF ordering (no spin canting). For the SN state, the enhancement of Gilbert damping for $\theta=\pi/2$ is twice that of $\theta=0$. In general, $\delta\alpha_{\rm G}$ is proportional to $1+\sin^{2}\theta$. This dependence arises from the fact that $\delta\alpha_{\rm G}$ is a sum of the two contributions from two types of Schwinger bosons; the contribution of the $z^{\prime}$-component Schwinger bosons is independent of $\theta$, while that of the $x^{\prime}$-component Schwinger bosons is proportional to $\sin^{2}\theta$ (see also Eqs. (59)-(61)). In contrast, there is no angle dependence for the antiferromagnetic insulator as shown in Fig. 3 (b). This corresponds to the fact that, in the absence of a magnetic field, only a collinear antiferromagnetic order is established, exhibiting rotational symmetry around the $x^{\prime}$-axis.
IV.2 Finite magnetic field
Figure 4: The increase of the Gilbert damping as a function of the FMR frequency $\omega_{\bm{0}}$ at $\theta=0$ for (a) the SN insulator and (b) the CAF insulator. The parameters are set as $J_{1}=0$, $J_{2}=-{\cal J}$, and $h=3{\cal J}$. Red arrows indicate the energy gap.
@@ IV.2 Finite magnetic field
Figure 5: The increase of the Gilbert damping for the SN case is shown as a function of the FMR frequency $\omega_{\bm{0}}$ for $J_{1}=0$ and $J_{2}=-{\cal J}$. (a) The magnetic-field dependence at $\theta=0$. The three curves correspond to $h={\cal J}$, $3{\cal J}$, and $5{\cal J}$, respectively. (b) The angle dependence at $h=3{\cal J}$. The red, blue, and green solid lines correspond to the contribution of $C_{z}$ for $\theta=0$, $\theta=\pi/2$, and $\theta=\pi$, respectively. The dashed line represents the contribution of $C_{\perp}$ at $\theta=\pi/2$. We note that $C_{\perp}(\theta=0)=C_{\perp}(\theta=\pi)=0$.
Fig. 4 shows $\delta\alpha_{\rm G}/\delta\alpha_{{\rm G},0}$ as a function of the FMR frequency $\omega_{\bm{0}}$ in the presence of an external magnetic field at $\theta=0$. For the SN insulator, $\delta\alpha_{\rm G}$ becomes zero at frequencies below the energy gap (indicated by the red arrow), as shown in Fig. 4 (a). However, for the CAF insulator, a finite value remains, as shown in Fig. 4 (b). This is because, at $\theta=0$, only the gapped bosons contribute to the spin pumping in the SN case, whereas both gapped and gapless bosons contribute in the CAF case (see also Eqs. (59)-(61) and Eqs. (64)-(65)).
@@
To clarify the magnetic field dependence in the SN case, we show $\delta\alpha_{\rm G}/\delta\alpha_{{\rm G},0}$ as a function of the FMR frequency $\omega_{\bm{0}}$ in the presence of an external magnetic field in Fig. 5. As the magnetic field increases, $\delta\alpha_{{\rm G}}(\omega_{\bm{0}})$ grows because the net magnetization in the spin-nematic insulator increases and the peak becomes sharper reflecting changes in the dispersion relation of the Schwinger bosons (see also Fig. 6(d)-(f)). Fig. 5 (b) shows the increase of the Gilbert damping at $\theta=0$, $\theta=\pi/2$, $\theta=\pi$ with the magnetic field fixed at $h=3{\cal J}$, where the contributions of the two kinds of Schwinger bosons are plotted separately. We note that the contribution from the perpendicular Schwinger bosons vanishes for $\theta=0,\pi$. The peak at $\omega_{\bm{0}}=2.9{\cal J}$ is the contribution of the $z^{\prime}$-component Schwinger bosons, and the peak at $\omega_{\bm{0}}=3.2{\cal J}$ is that of the perpendicular component of the Schwinger bosons. As seen in the figure, the two components of the bosons have different dependencies on $\theta$.
Figure 6: (a)-(c) The color gradient in the three-dimensional plot represents the increase of the Gilbert damping, $\delta\alpha_{\rm G}$, of the SN/FI junction as a function of the FMR frequency $\omega_{\bm{0}}$ and $\theta$ for $h={\cal J}$, $3{\cal J}$, and $5{\cal J}$, respectively. (d)-(f) The corresponding dispersion of the two Schwinger bosons of $\varepsilon_{{\bm{k}}z}$ and $\varepsilon_{{\bm{k}}\perp}$. (g)-(i) The color gradient in the three-dimensional plot represents the increase of the Gilbert damping, $\delta\alpha_{\rm G}$, of the CAF/FI junction for $h={\cal J}$, $3{\cal J}$, and $5{\cal J}$, respectively. (j)-(l) The corresponding dispersion of the two magnon modes, $\varepsilon_{{\bm{k}}\alpha}$ and $\varepsilon_{{\bm{k}}\beta}$.
@@
To compare the dependence on the magnetic-field orientation between the SN/FI and CAF/FI junctions, we show the increase of the Gilbert damping as a function of $\omega_{\bm{0}}$ and $\theta$ for $h={\cal J}$, $3{\cal J}$, and $5{\cal J}$ in Fig. 6(a)-(c) and (g)-(i), where the first column (third column) corresponds to the SN/FI (CAF/FI) junction. For both junctions, the peak height at a finite frequency varies as the angle $\theta$ changes. The shape and position of the peak of $\delta\alpha_{\rm G}$ are determined by the corresponding dispersion of the two bosons, as shown in Fig. 6(d)-(f) for the SN case and (j)-(l) for the CAF case. As expected from the form of Eq. (59) and Eq. (64), the peak frequency of $\delta\alpha_{\rm G}$ corresponds to the energy of the flat part in the dispersion of the two bosons in both cases. The low-frequency shoulder structure of $\delta\alpha_{\rm G}$ reflects the low-energy states of the Schwinger bosons, though its weight is much weaker than that of the main peak.
@@
Let us examine the detailed features of the SN/FI junction under a finite magnetic field. In Fig. 6(a)-(c), we show three-dimensional plots of $\delta\alpha_{\rm G}$ as a function of the resonant frequency $\omega_{\bm{0}}$ and the angle $\theta$ for the SN/FI junction. For comparison, we also show the same plots for the CAF/FI junction in Fig. 6(g)-(i). For $h={\cal J}$ (Fig. 6(a)), $\delta\alpha_{\rm G}$ has a single peak around $\omega_{\bm{0}}\simeq 3{\cal J}$ and this peak becomes most significant near $\theta=\pi/2,3\pi/2$. For a higher magnetic field of $h=3{\cal J}$ (Fig. 6(b)), a secondary peak appears in the high-energy side with a different $\theta$-dependence; the lower-energy peak around $\omega_{\bm{0}}=2.9{\cal J}$ becomes significant near $\theta=0$, while the high-energy peak around $\omega_{\bm{0}}=3.2{\cal J}$ appears around $\theta=\pi/2,3\pi/2$. The positions of the double peaks correspond to the flat bands, whose energies are different for the two types of bosons (see Fig. 6(e)), while their weights have different dependencies on $\theta$, as seen also in Eqs. (59)-(61). For the highest magnetic field of $h=5{\cal J}$ (Fig. 6(c)), the contribution of the perpendicular component of the Schwinger bosons almost vanishes, while that of the $z^{\prime}$-component bosons becomes dominant. As a result, $\delta\alpha_{\rm G}$ has only a single peak due to the flat part in the dispersion of the $z^{\prime}$ Schwinger bosons. For the CAF/FI junction (Fig. 6(g)-(i)), there always appears a single peak regardless of the magnetic field strength, whose height has a maximum value at $\theta=0,2\pi$. It is remarkable that no peaks are observed near $\theta=\pi/2,3\pi/2$ as in the SN/FI junction.
@@ V Summary
In this study, we examined spin pumping into a spin-nematic (SN) state in a junction system composed of a ferromagnetic insulator and a spin-nematic insulator. To clarify its qualitative features, we employed a spin-1 bilinear-biquadratic model to realize the spin-nematic phase and calculated the increase of the Gilbert damping using the mean-field theory based on the Schwinger boson formalism. To clarify the feature of the SN state, we compared the obtained results for the model of the canted antiferromagnetic insulator (CAF) because both the SN and CAF states have two bosonic modes and do not conserve a spin component in a specific direction due to the U(1) symmetry breaking. Although these two models lead to similar results at first glance, there are clear differences between them. (i) For zero magnetic field, $\delta\alpha_{\rm G}$ depends on the spin angle of the FI, $\theta$, for the SN/FI junction, while it is independent of $\theta$ for the CAF/FI junction (see Fig. 3). (ii) For a finite magnetic field, $\delta\alpha_{\rm G}$ vanishes below a specific frequency for the SN/FI junction, while it continues to be finite for the CAF/FI junction (see Fig. 4). (iii) For a finite magnetic field, the peaks in $\delta\alpha_{\rm G}$ have a different $\theta$-dependence between the SN/FI and CAF/FI junctions (see Fig. 6); For the CAF/FI junction, there appears a single peak with $\cos\theta$ type dependence, while for the SN/FI junction another peak with $\cos 2\theta$ type dependence coexists when the magnetic field is not too strong. These findings are also consistent with analytic results in Sec. III.
Our study demonstrates that spin pumping, a well-established technique in spintronics, provides a promising route to identify the spin-nematic (SN) phase. By tuning the microwave frequency, this approach allows systematic control of the relevant energy scale, while altering the orientation of the applied magnetic field enables clear separation of multiple magnetic excitations. Such features are particularly advantageous for probing the unique excitation spectra characteristic of the SN phase.
Finally, we comment on the experimental realization. The exchange interactions of the candidate materials, ${\mathrm{LiCuVO}}_{4}$ and ${\mathrm{BaCdVO}}(\mathrm{PO}4)_{2}$, for spin nematic states are of the order of $0.5$-$5\,{\rm meV}$, which corresponds to the frequency of $0.1$-$1\,{\rm THz}$. This frequency scale is much larger than that used in the usual FMR experiment. Therefore, to observe the main peaks at the frequency scale of ${\cal J}$, we need to use spin systems with smaller exchange interactions or to apply high-speed THz spin pumping technique [73, 74]. On the other hand, the absence of the increase of the Gilbert damping at low frequencies (see Fig. 4(a)) may be observed using the usual FMR experiment. Study on detailed conditions for observation in realistic candidates of the SN phase is left as future problems.
@@ Acknowledgements
The authors thank M. Matsuo and M. Sato for their helpful discussions. T. I. is supported by the International Graduate Program of Innovation for Intelligent World (IIW) of the University of Tokyo and the International Program of the Institute for Solid State Physics, the University of Tokyo and acknowledges the support and hospitality of the IRTG Fluctuations and Nonlinearities of SFB 1432 during his stay at University of Konstanz. T. K. acknowledges the support of the Japan Society for the Promotion of Science (JSPS KAKENHI Grant No. JP24K06951). W. B. acknowledges financial support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Project-ID 425217212 - SFB 1432.
@@ Appendix A Detailed calculation of the spin-spin correlation function
$S^{+^{\prime}}_{\bm{k}}$ $=i\sqrt{M}\,\bigl{[}(C^{-}_{\mu}c^{z}_{{\bm{k}}}-C^{+}_{\mu}s^{z}_{{\bm{k}}})\alpha_{{\bm{k}}z}^{\dagger}+(C^{-}_{\mu}s^{z}_{{\bm{k}}}-C^{+}_{\mu}c^{z}_{{\bm{k}}})\alpha_{-{\bm{k}}z}\bigr{]},$ (68)
@@
$\chi^{z^{\prime}z^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)$ $=M\cos^{2}\mu(c_{\bm{k}}^{\perp}-s_{\bm{k}}^{\perp})^{2}\bigl{[}G_{{\bm{k}}\perp}(\tau)+G_{{\bm{k}}\perp}(-\tau)\bigr{]},$ (73)
@@
Using the diagonalized Hamiltonian, Eq. (23), the Fourier transformation of this propagator is calculated as
@@
Appendix B Detailed calculation of the spin-spin correlation function in canted antiferromagnetic insulator
In this appendix, we provide a detailed calculation of the spin-spin correlation functions for the CAF insulator. We first rewrite the spin operators in the $(x^{\prime},y^{\prime},z^{\prime})$ coordinate system using the magnon operators, following Eqs. (28) and (29). We then apply the Bogoliubov transformation given in Eqs. (30), (35) and (36). As a result, the spin operators are calculated as
$S^{+^{\prime}}_{A,\bm{k}}$ $=\frac{\sqrt{S}}{2}(\cos\phi+1)(\cosh\psi_{\bm{k}}^{\alpha}\alpha_{\bm{k}}+\sinh\psi_{\bm{k}}^{\alpha}\alpha^{\dagger}_{-\bm{k}}+\cosh\psi_{\bm{k}}^{\beta}\beta_{-\bm{k}}+\sinh\psi_{\bm{k}}^{\beta}\beta^{\dagger}_{\bm{k}})$
$\hskip 14.22636pt+\frac{\sqrt{S}}{2}(\cos\phi-1)(\sinh\psi_{\bm{k}}^{\alpha}\alpha_{\bm{k}}+\cosh\psi_{\bm{k}}^{\alpha}\alpha^{\dagger}_{-\bm{k}}+\sinh\psi_{\bm{k}}^{\beta}\beta_{-\bm{k}}+\cosh\psi_{\bm{k}}^{\beta}\beta^{\dagger}_{\bm{k}})-S\sin\phi,$ (76)
$S^{z^{\prime}}_{A,\bm{k}}$ $=\frac{\sqrt{S}\sin\phi}{2}(\cosh\psi_{\bm{k}}^{\alpha}\alpha_{\bm{k}}+\sinh\psi_{\bm{k}}^{\alpha}\alpha^{\dagger}_{-\bm{k}}+\cosh\psi_{\bm{k}}^{\beta}\beta_{-\bm{k}}+\sinh\psi_{\bm{k}}^{\beta}\beta^{\dagger}_{\bm{k}})$
$\hskip 14.22636pt+\frac{\sqrt{S}\sin\phi}{2}(\sinh\psi_{\bm{k}}^{\alpha}\alpha_{\bm{k}}+\cosh\psi_{\bm{k}}^{\alpha}\alpha^{\dagger}_{-\bm{k}}+\sinh\psi_{\bm{k}}^{\beta}\beta_{-\bm{k}}+\cosh\psi_{\bm{k}}^{\beta}\beta^{\dagger}_{\bm{k}})+S\cos\phi,$ (77)
$S^{+^{\prime}}_{B,\bm{k}}$ $=\frac{\sqrt{S}}{2}(\cos\phi+1)(-\cosh\psi_{\bm{k}}^{\alpha}\alpha_{-\bm{k}}-\sinh\psi_{\bm{k}}^{\alpha}\alpha^{\dagger}_{\bm{k}}+\cosh\psi_{\bm{k}}^{\beta}\beta_{\bm{k}}+\sinh\psi_{\bm{k}}^{\beta}\beta^{\dagger}_{-\bm{k}})$
$\hskip 14.22636pt+\frac{\sqrt{S}}{2}(\cos\phi-1)(-\sinh\psi_{\bm{k}}^{\alpha}\alpha_{-\bm{k}}-\cosh\psi_{\bm{k}}^{\alpha}\alpha^{\dagger}_{\bm{k}}+\sinh\psi_{\bm{k}}^{\beta}\beta_{\bm{k}}+\cosh\psi_{\bm{k}}^{\beta}\beta^{\dagger}_{-\bm{k}})+S\sin\phi,$ (78)
$S^{z^{\prime}}_{B,\bm{k}}$ $=-\frac{\sqrt{S}\sin\phi}{2}(-\cosh\psi_{\bm{k}}^{\alpha}\alpha_{-\bm{k}}-\sinh\psi_{\bm{k}}^{\alpha}\alpha^{\dagger}_{\bm{k}}+\cosh\psi_{\bm{k}}^{\beta}\beta_{\bm{k}}+\sinh\psi_{\bm{k}}^{\beta}\beta^{\dagger}_{-\bm{k}})$
$\hskip 14.22636pt-\frac{\sqrt{S}\sin\phi}{2}(-\sinh\psi_{\bm{k}}^{\alpha}\alpha_{-\bm{k}}-\cosh\psi_{\bm{k}}^{\alpha}\alpha^{\dagger}_{\bm{k}}+\sinh\psi_{\bm{k}}^{\beta}\beta_{\bm{k}}+\cosh\psi_{\bm{k}}^{\beta}\beta^{\dagger}_{-\bm{k}})+S\cos\phi,$ (79)
and $S_{\nu,{\bm{k}}}^{-^{\prime}}=(S_{\nu,-{\bm{k}}}^{+^{\prime}})^{\dagger}$ ($\nu={\rm A},{\rm B}$). Using these operators, the spin-spin correlation functions are calculated as
$\chi^{z^{\prime}z^{\prime}}_{{\bm{k}},-{\bm{k}}}(\tau)=$ $\frac{S\sin^{2}\phi}{2}\Bigl{[}(K^{+}_{\bm{k}\alpha})^{2}\Bigl{\{}G_{\bm{k}\alpha}(\tau)+G_{\bm{k}\alpha}(-\tau)\Bigr{\}}+(K^{+}_{\bm{k}\beta})^{2}\Bigl{\{}G_{\bm{k}\beta}(\tau)+G_{\bm{k}\beta}(-\tau)\Bigr{\}}\Bigr{]},$ (80)
$\chi^{+^{\prime}+^{\prime}}_{{\bm{k}},{-\bm{k}}}(\tau)=$ $\frac{S}{2}\Bigl{\{}(K^{+}_{\bm{k}\alpha})^{2}\cos^{2}\phi-(K^{-}_{\bm{k}\alpha})^{2}\Bigr{\}}\Bigl{[}G_{\bm{k}\alpha}(\tau)+G_{\bm{k}\alpha}(-\tau)\Bigr{]}$
$\hskip 5.69054pt+\frac{S}{2}\Bigl{\{}(K^{+}_{\bm{k}\beta})^{2}\cos^{2}\phi-(K^{-}_{\bm{k}\beta})^{2}\Bigr{\}}\Bigl{[}G_{\bm{k}\beta}(\tau)+G_{\bm{k}\beta}(-\tau)\Bigr{]},$ (81)
$\chi^{+^{\prime}-^{\prime}}_{{\bm{k}},{-\bm{k}}}(\tau)=$ $\frac{S}{2}\Bigl{[}\Bigl{\{}K^{+}_{\bm{k}\alpha}\cos\phi+K^{-}_{\bm{k}\alpha}\Bigr{\}}^{2}G_{\bm{k}\alpha}(\tau)+\Bigl{\{}K^{+}_{\bm{k}\beta}\cos\phi+K^{-}_{\bm{k}\beta}\Bigr{\}}^{2}G_{\bm{k}\beta}(\tau)\Bigr{]}$
$\hskip 5.69054pt+\frac{S}{2}\Bigl{[}\Bigl{\{}K^{+}_{\bm{k}\alpha}\cos\phi-K^{-}_{\bm{k}\alpha}\Bigr{\}}^{2}G_{\bm{k}\alpha}(-\tau)+\Bigl{\{}K^{+}_{\bm{k}\beta}\cos\phi-K^{-}_{\bm{k}\beta}\Bigr{\}}^{2}G_{\bm{k}\beta}(-\tau)\Bigr{]},$ (82)
$\chi^{-^{\prime}+^{\prime}}_{{\bm{k}},{-\bm{k}}}(\tau)=$ $\frac{S}{2}\Bigl{[}\Bigl{\{}K^{+}_{\bm{k}\alpha}\cos\phi-K^{-}_{\bm{k}\alpha}\Bigr{\}}^{2}G_{\bm{k}\alpha}(\tau)+\Bigl{\{}K^{+}_{\bm{k}\beta}\cos\phi-K^{-}_{\bm{k}\beta}\Bigr{\}}^{2}G_{\bm{k}\beta}(\tau)\Bigr{]}$
$\hskip 5.69054pt+\frac{S}{2}\Bigl{[}\Bigl{\{}K^{+}_{\bm{k}\alpha}\cos\phi+K^{-}_{\bm{k}\alpha}\Bigr{\}}^{2}G_{\bm{k}\alpha}(-\tau)+\Bigl{\{}K^{+}_{\bm{k}\beta}\cos\phi+K^{-}_{\bm{k}\beta}\Bigr{\}}^{2}G_{\bm{k}\beta}(-\tau)\Bigr{]},$ (83)
and $\chi^{-^{\prime}-^{\prime}}_{{\bm{k}},{-\bm{k}}}(\tau)=\chi^{+^{\prime}+^{\prime}}_{{\bm{k}},{-\bm{k}}}(\tau)$, where $\cosh\psi_{\bm{k}}^{\nu}\pm\sinh\psi_{\bm{k}}^{\nu}=K^{\pm}_{{\bm{k}}\nu}$ and $G_{{\bm{k}}\nu}(\tau)$ are propagators of the magnons defined by
$G_{{\bm{k}}\nu}(\tau)=-\frac{1}{\hbar}\langle\nu_{{\bm{k}}}(\tau)\nu^{\dagger}_{{\bm{k}}}(0)\rangle.$ (84)
Using the diagonalized Hamiltonian, Eq. (39), the Fourier transformation of this propagator is calculated as
[…]